OpenStudy (anonymous):

What is the solution for x in the proportion 12 / 22 = x / 99.? Describe your steps. My Answer: 12/22 = x/99 (multiplying by 99) (12 * 99)/22 = x x = (12 * 9)/2 (dividing numerator and denominator by 11 in right hand side x = 6 * 9 x = 54 Is this right? did I do anything wrong or am I good to go?

OpenStudy (anonymous):

this equation is easy just you can applay the cross product directly and you find . we have 12/22=x/99 implies that 12*99=x*22 as a result x=(12*99)/22 calculate you faind the value

OpenStudy (anonymous):

right multiply 12*99=22*x 12*99/22=54

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

NP.

OpenStudy (anonymous):

you are welcom

OpenStudy (anonymous):

:)